OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x2 + ___x + 49 help I will award medals

OpenStudy (whpalmer4):

If it's a perfect square, the leading term must be the square of the leading term of the product binomial. So \(4x^2 = a^2, a =\)

OpenStudy (whpalmer4):

Similarly, the constant term must be the square of the constant term of the product binomial, so \(49 = b^2, b =\)

OpenStudy (whpalmer4):

And \[(a+b)^2 = (a+b)(a+b) = a(a+b) + b(a+b) = a^2+ab + ab + b^2 \]\[= a^2+2ab + b^2\]So the middle term (aka the missing term) = \(2 a b\) where \(a,b\) are the values you found earlier...

OpenStudy (whpalmer4):

(don't forget to remove the \(x\) when typing in the answer!\)

OpenStudy (anonymous):

4x^2+2x+49?

OpenStudy (whpalmer4):

If \[a^2 = 4x^2\]take square root of both sides\[\sqrt{a^2}=\sqrt{4x^2}\]\[a = \sqrt{4}*\sqrt{x^2} = 2x\]right?

OpenStudy (whpalmer4):

if \[b^2=49\]take square root of both sides\[\sqrt{b^2} = \sqrt{49}\]\[b = \sqrt{49}\]\[b=7\]right?

OpenStudy (whpalmer4):

missing term is then\[2ab = 2(2x)(7) =\]

OpenStudy (anonymous):

So the answer should be 4x^2+2x+49

OpenStudy (whpalmer4):

No! what is \(2(2x)(7)\)

OpenStudy (anonymous):

What?

OpenStudy (whpalmer4):

what is \(2*2x*7\)

OpenStudy (anonymous):

28x

OpenStudy (whpalmer4):

Right! so the missing term is \(28x\). Here, let me show you. We have a perfect square trinomial, which means we have \((a+b)^2 = 4x^2 + <something> + 49\) and we established that \(a = 2x,~b = 7\) That gives us \[(2x+7)(2x+7) = 2x(2x+7)+7(2x+7) = 4x^2 + 14x + 14x + 49 \]\[=4x^2 + 28x + 49\] Therefore \(28x\) is the missing term, and you enter "28" as your answer (because they don't want you to include the variable, \(x\).

OpenStudy (anonymous):

Thanks

OpenStudy (whpalmer4):

when you square a binomial, the middle term is always 2 * first term * second term. We just had to puzzle out what the first and second terms must have been...